# Entanglement dynamics in intensity-dependent double Jaynes-Cummings model for squeezed coherent thermal states

### v4

@@ 4 Effects of single photon exchange interaction on the entanglement dynamics of the atom-field system
Figure 7 shows the effects of photon exchange interaction for SCTS in DJCM. The maximum value of of $N(t)$ for field-field subsystem increases by almost three times as a result of photon exchange interaction. For $\kappa=0.1$, which is a low value compared to the atom-field coupling constant $\lambda$. For this low value of $\kappa$, the lengths of ESD for atom-atom entanglement increase and other smaller peaks disappear. The atom A-field a entanglement comes down significantly and at some places, $N(t)$ almost becomes zero. In the case of the atom A-field b entanglement, we observed that the ESDs are removed from the dynamics of $N(t)$. For $\kappa=1.0$, the length of ESD of $C(t)$ becomes larger as small peaks disappear. $N(t)$ for field-field entanglement increases very sharply and then comes down again. Its maximum value also decreases as compared to the case when $\kappa=0.1$. In atom A-field a subsystem, ESD occurs as a result of increase in $\kappa$. No, considerable change is noticed in atom A-field a entanglement. As $\kappa$ is increased further to $5$, a drastic change is observed in the entanglement dynamics of $C(t)$. We see that all the ESDs are removed from the dynamics and its value also increases significantly at every instant of time. In field-field entanglement, we notice that the maximum value of $N(t)$ increases and the dynamics becomes very oscillatory. $N(t)$ for atom A-field a and atom A-field b shows ESD initially and after that it stays nonzero with very low amplitude. When $\kappa>>\lambda$, i.e., for $\kappa=10$, $C(t)$ increases more and remains high. We observe interesting pattern in the dynamics of $N(t)$ for field-field entanglement. In this case wave packets of entanglement are formed and amplitude of the wave packets increase significantly. $N(t)$ for atom A-field a and atom A-field b entanglements becomes zero. From this analysis we conclude that photon exchange interaction creates entanglement in the atom-atom and field-field subsystems. However, in the case of IDDJCM, we observe that $C(t)$ increases and $N(t)$ for field-field entanglement decreases; however, in the case of DJCM both $C(t)$ and $N(t)$ for the field-field entanglement increase. In both IDDJCM and DJCM, atom A-field a and atom A-field b entanglements become zero with increasing $\kappa$.
5 Effects of Kerr-nonlinearity on the entanglement dynamics in IDDJCM
@@
The effects of $\chi$ on entanglement dynamics of different subsystems for SCTS in DJCM are presented in Fig. 9. For $\chi=0.1$, ESD in $C(t)$ increases considerably since the smaller peaks disappear from the dynamics (Fig. 9(a), blue curve). For the other subsystems, the dynamics of $N(t)$ does not change considerably for small values of $\chi$. As $\chi$ is increased to $0.3$, the amplitudes of $C(t)$ and $N(t)$ increase; however, ESDs are not removed from the dynamics. For other entanglements, $N(t)$ for atom A-field a and field a-field b decrease prominently with the increasing nonlinearity in the system. However, $N(t)$ for the atom A-field b subsystem increases and ESDs are removed from the dynamics. On further increasing the value of $\chi$ to $0.7$ and $1.0$, $C(t)$ increases very significantly with greater amplitude and ESDs are removed from the dynamics. Similar behaviour is observed for atom A-field b entanglement dynamics. On the other hand, for atom A-field a and field-field entanglement, $N(t)$ decreases prominently with these increased values. This shows that with the addition of $\chi$ to the system, the entanglement is transferred in the atom-atom and atom A-field b subsystems from the atom A-field b and field a-field b subsystems.
6 Effects of spin-spin Ising interaction between the two atoms on entanglement dynamics
and
where $J_{z}$ is the coupling strength between the two atoms. $J_{z}$ has the unit of energy. Recently people have studied the Ising-type interaction for different configurations. Ghoshal et. al., have investigated the entanglement dynamics of the quenched disordered double Jayens-Cummings model in the presence of Ising-type interaction in the Hamiltonian[31]. In another work, Pandit et. al., have also studied the effects of Ising-type interaction on the entanglement dynamics in the DJCM [28]. Ref.[33] has analyzed the effects of Ising-type interaction on entanglement in DJCM for vacuum, coherent and thermal states of radiation fields. Sadiek et. al.,[80] have investigated the time evolution and asymptotic behaviour of entanglement with including the dipole-dipole and Ising-type interactions in the Hamiltonian.
The effects of Ising interaction on entanglement for IDDJCM are shown in Fig. 10. The plots show very different effects as are observed in DJCM for SCS and G-L states. From Fig. 10(a) and 10(b), it can be seen that for $J_{z}=0.1$ and $0.3$, we cannot distinguish the plots from the plots with $J_{z}=0$ for $C(t)$ and $N(t)$ (see Figs. 1 and 10). However, for atom A-field b entanglement, the amplitude of $N(t)$ comes down and ESDs are also present in the dynamics. In the case of field a-field b entanglement, for the above mentioned values of $J_{z}$, we observe that ESDs are removed from the dynamics and amplitude increases. For $J_{z}=0.7$, small peaks appear along with the large peaks in the dynamics of $C(t)$ (red curve) and also the amplitude decreases. For atom A-field a subsystem, the amplitude of $N(t)$ comes down a little and ESDs are also removed in due course of time. In the case of atom A-field b, we do not see any noticeable change in the dynamics. However, for the field a-field b subsystem, the amplitude increases further, and no ESD is observed. For $J_{z}=1.0$, small peaks in the dynamics of $C(t)$ appear earlier in the dynamics and its height increases with time, while the initial peak decreases. As a result, the duration of the ESDs gets smaller. For other subsystems, $N(t)$ follows the dynamics similar to that corresponding to $J_{z}=0.7$, viz., with decreased peak height for atom A-field a entanglement while increased peak height for atom A-field b and field a-field b subsystems.
7 Effects of detuning on the entanglement dynamics for SCTS in and IDDJCM and DJCM
@@
To study the effects of detuning $\bar{n}_{c}=2,\bar{n}_{s}=1.0$ and $\bar{n}_{th}=0.1$ are chosen while the values of $\Delta$ are taken to be $2,5,10$. The effects of detuning on entanglements for SCTS in IDDJCM is plotted in Fig. 12. From the plots it is very much noticeable that $\Delta$ plays a significant role in removing ESDs from the system. For $\Delta=2$, length of ESDs get decreased in the dynamics of $C(t)$ (blue curve in Fig. 12(a)) by introducing new smaller peaks in the dynamics. Another important fact is that the detuning decreases the peak heights of the entanglement pulses significantly with increasing time; however the periodic nature of $C(t)$ is present to some extent. For atom A-field a entanglement, we see that the addition of $\Delta$ in the system totally destroys the periodic nature of the dynamics of $N(t)$. The amplitude of $N(t)$ is increased and all the ESDs are removed from the dynamics. The periodic pulse like nature of $N(t)$ for atom A-field b is also get destroyed by the addition of $\Delta$ in the system. For this smaller value of $\Delta$, length of the ESDs get smaller however these are not removed from the dynamics. In the case of field a-field b subsystem it is observed that the addition of $\Delta=2$ removes all the ESDs from the dynamics of $N(t)$. Although, the periodic nature of $N(t)$ is present in the dynamics, the form of the periodic nature of is distorted significantly.
@@
For $\Delta=5$, the height of the peaks of $C(t)$ comes down further with smaller peaks appearing in some places of the dynamics. This makes the length of ESDs larger at some places and smaller at some places. Also the entanglement pulses are destroyed with increasing $\Delta$. $N(t)$ for atom A-field a does not change considerably, while for atom A-field b the amplitude of $N(t)$ increases and some of the ESDs get removed from the dynamics. For field a-field b subsystems, it is observed that the amplitude of $N(t)$ becomes almost half of the initial peak and the periodicity of the dynamics is totally lost.
@@
As $\Delta$ is increased to $\Delta=10$, the length of initial ESD increases as a result of disappearance of other peaks from the dynamics. The dynamics becomes oscillatory for a short time and again ESDs appear. So, we can see that detuning is not much effective to remove the ESDs from the dynamics of $C(t)$ and increases the value of atom-atom entanglement which is observed for SCS and G-L states.
@@
The effects of $\Delta$ for SCTS in DJCM is depicted in Fig. 13. For $\Delta=2$, initial ESD is removed from the dynamics of $C(t)$, however the peak after the second long ESD is lost (see black curve in Fig. 3(a) and blue curve in Fig. 13(a)), which makes the second ESD even longer. For the all the other subsystems, the addition of $\Delta$ decreases the amplitude $N(t)$. For the higher values of $\Delta=5,10$, though the amplitude of $C(t)$ increases more initially, the length of later ESDs increases significantly. In the case of other subsystems, $N(t)$ is decreased further which makes ESD appear initially in field-field entanglement. From these plots it is evident that presence of thermal photons bar entanglement transfer to atom-atom subsystem from other subsystems.
8 Effects of dipole-dipole interaction between the two atoms
In this section, the effects of dipole-dipole interaction on entanglement dynamics are discussed. The intensity-dependent double JC Hamiltonian with dipole-dipole interaction can be written as
Much interest have been shown in studying the effects of the dipole-dipole interaction on the dynamics of the atom-field system for a long time. In particular, the authors in [81], have discussed a detailed dynamics of the two-atom system where two atoms are interacting with each other via dipole-dipole coupling. A remarkable feature which comes out of this study is that more series of revivals appear in the dynamics that are like the one-atom case, partial and overlapping. They have also found that quantum collapse is no longer Gaussian and now depends on the dipole-dipole interaction parameter. In Ref.[82], Evseev et. al., have investigated entanglement dynamics between two qubits in a non-resonant DJCM taking into account the direct dipole-dipole interaction between the qubits. The results show that the dipole-dipole parameter has great impact on the entanglement dynamics. They have also shown that the presence of large $g_{d}$ leads to a stabilization of entanglement for all Bell-types initial qubit states and different couplings and detunings. In another paper[83] , the authors have studied the nonlinear version of the Jaynes-Cummings model for two identical two-level atoms for Ising-like and dipole-dipole interaction between the atoms. The have found that when the ratio of $\frac{g_{d}-J_{z}}{\lambda}>>1$, some significant results are observed. The effects of dipole-dipole interaction on the energy levels of an one dimensional system using Jaynes-Cummings Hubbard model is discussed in Ref.[84].
The effects of the dipole-dipole interaction on entanglement dynamics for the IDDJCM is represented in Fig. 14. The effects for $g_{d}=0.1$ and $g_{d}=1.0$ are almost similar as we observe in photon exchange interaction for $\kappa=0.1$ and $1.0$. For $g_{d}=5$, we observe that field-field entanglement is different from the dynamics for $\kappa=5$. In this case no oscillation is found in the dynamics though its amplitude increases significantly. The behaviour of other entanglements are quite similar. When $g_{d}$ is increased to $10$, all the ESDs are removed from the dynamics of $C(t)$. Also the amplitude of oscillations are increased; however the oscillations die after sometime. For the field-field entanglement, it is noticed that the downward fall of the dynamics stops. The atom-field entanglements become zero for this high value of dipole-dipole interaction.
The effects of dipole-dipole interaction for DJCM with SCTS is plotted in Fig. 15. From Fig. 15(a), it can be seen that all the dynamics almost remain unchanged for $g_{d}=0.1$. Only field-field entanglement rises up for larger time making ESDs disappear. For $g_{d}=1.0$, peaks of $C(t)$ come down remarkably but length of ESDs remains same. The field-field entanglement increases more and all the ESDs are removed from the dynamics. The amplitude of $N(t)$ for atom A-field a comes down however, $N(t)$ for the atom A-field b does not change considerably. For $g_{d}=5$, it is observed that new peaks start to appear in the dynamics of $C(t)$ and all the ESDs are removed the dynamics for $g_{d}=10$. The field-field entanglement remains high as before. However, the atom-field entanglements almost becomes zero for $g_{d}=10$.
@@ 9 Conclusion
In this work, we have investigated the entanglement dynamics of the atom-atom, atom-field and field-field subsystems with squeezed coherent thermal states in intensity-dependent double Jaynes-Cummings model (IDDJCM) and standard double Jaynes-Cummings model (DJCM). It is noticed that the atom-atom and field-field entanglements almost remain unchanged with increasing number of squeezed photons while keeping coherent and thermal photons constant for IDDJCM. An array of entanglement pulses is formed with almost same ESDs and same peak amplitude. However, the atom-field entanglements are affected noticeably. For DJCM, all the dynamics change with increasing squeezed photons and length of ESDs also increases for atom-atom and atom A-field b subsystems.
The addition of photon exchange interaction between the cavities removes ESDs from field-field entanglement dynamics and from the atom-atom for large values of $\kappa$ for both IDDJCM and DJCM. The atom-field entanglements become zero for large values of $\kappa$. In the case of DJCM, formation of wave packets is observed for field-field subsystem. Kerr-nonlinearity effectively removes ESDs from the all the entanglement dynamics. For IDDJCM, $C(t)$ and $N(t)$ for all the subsystems increase for higher values of $\chi$. In the case of DJCM also, $C(t)$ and $N(t)$ for atom-field subsystems increase for higher values of $\chi$, however, field-field entanglement decreases after a critical value of $\chi$.
Ising interaction is effective to remove ESDs from field-field entanglement for IDDJCM; however, it is unable to remove ESDs from atom-atom and atom-field entanglements. For DJCM, the dynamics of $C(t)$ and $N(t)$ are not affected considerably by the addition of Ising interaction in the system. Detuning tries to remove ESDs from all the subsystems. ESDs are removed from field-field entanglement, however its amplitude comes down with increasing value of $\Delta$. Also, for higher values of detuning, the $C(t)$ increases more and other entanglements decreases which is a implications of entanglement transfer from other subsystems to atom-atom subsystems. Like photon exchange interaction, dipole-dipole interaction removes ESDs from the dynamics of $C(t)$ and $N(t)$ of field-field subsystem for high value of $g_{d}$ for both IDDJCM and DJCM. However, in this case, formation of wave packets is not observed.

### v5

@@ 4 Effects of single photon exchange interaction on the entanglement dynamics
Figure 7 shows the effects of photon exchange interaction for SCTS in DJCM. The maximum value of of $N(t)$ for field-field subsystem increases by almost three times as a result of photon exchange interaction. For $\kappa=0.1$, which is a low value compared to the atom-field coupling constant $\lambda$. For this low value of $\kappa$, the lengths of ESDs for atom-atom entanglement increase and other smaller peaks disappear. The atom A-field a entanglement comes down significantly and at some places, $N(t)$ almost becomes zero. In the case of the atom A-field b entanglement, we observe that the ESDs are removed from the dynamics of $N(t)$. For $\kappa=1.0$, the length of ESD of $C(t)$ becomes larger, as small peaks disappear. $N(t)$ for field-field entanglement increases very sharply and then comes down again. Its maximum value also decreases as compared to the case when $\kappa=0.1$. In atom A-field a subsystem, ESD occurs as a result of increase in $\kappa$. No appreciable change is noticed in atom A-field a entanglement. As $\kappa$ is increased further to $5$, a drastic change is observed in the entanglement dynamics of $C(t)$. We see that all the ESDs are removed from the dynamics and its value also increases significantly at every instant of time. In field-field entanglement, we notice that the maximum value of $N(t)$ increases and the dynamics becomes very oscillatory. $N(t)$ for atom A-field a and atom A-field b shows ESD initially and after that it stays nonzero with very low amplitude. When $\kappa>>\lambda$, i.e., for $\kappa=10$, $C(t)$ increases more and remains high. We observe interesting patterns in the dynamics of $N(t)$ for field-field entanglement. In this case, wave packets of entanglement are formed and amplitude of wave packets increase significantly. $N(t)$ for atom A-field a and atom A-field b entanglements become zero. From this, we conclude that the photon exchange interaction creates entanglement in the atom-atom and field-field subsystems. However, in the case of IDDJCM, we observe that $C(t)$ increases and $N(t)$ for field-field entanglement decreases. On the other hand, in the case of DJCM, both $C(t)$ and $N(t)$ for the field-field entanglement increase. In both IDDJCM and DJCM, atom A-field a and atom A-field b entanglements become zero with increasing $\kappa$.
5 Effects of dipole-dipole interaction between the two atoms on entanglement
In this section, the effects of dipole-dipole interaction on the entanglement dynamics are discussed. The intensity-dependent double JC Hamiltonian with dipole-dipole interaction can be written as
Much interest has gone into studying the effects of the dipole-dipole interaction on the dynamics of the atom-field system for a long time. In particular, the authors in [77], have discussed a detailed dynamics of the two-atom system where two atoms are interacting with each other via dipole-dipole coupling. A remarkable feature which comes out of this study is that more series of revivals appear in the dynamics that are like the one-atom case, partial and overlapping. It is also found that quantum collapse is no longer Gaussian and now depends on the dipole-dipole interaction parameter. In Ref.[78], Evseev et. al., have investigated the entanglement dynamics between two qubits in a non-resonant DJCM taking into account the direct dipole-dipole interaction between the qubits. The results show that the dipole-dipole parameter has great impact on the entanglement dynamics. They have also shown that the presence of large $g_{d}$ leads to a stabilization of entanglement for all Bell-type initial qubit states and different couplings and detunings. In another work [79] , the authors have studied the nonlinear version of the Jaynes-Cummings model for two identical two-level atoms for Ising-like and dipole-dipole interaction between the atoms. It is found that when the ratio $\frac{g_{d}-J_{z}}{\lambda}>>1$, certain significant results are observed. The effects of dipole-dipole interaction on the energy levels of an one-dimensional system using Jaynes-Cummings Hubbard model is discussed in Ref.[80].
The effects of the dipole-dipole interaction on the entanglement dynamics for the IDDJCM is represented in Fig. 8. The effects for $g_{d}=0.1$ and $g_{d}=1.0$ are almost similar, as we observe in photon exchange interaction for $\kappa=0.1$ and $1.0$. For $g_{d}=5$, we observe that field-field entanglement is different from the dynamics for $\kappa=5$. In this case no oscillation is found in the dynamics, though its amplitude increases significantly. The behaviour of other entanglements are quite similar. When $g_{d}$ is increased to $10$, all the ESDs are removed from the dynamics of $C(t)$. Also, the amplitude of oscillations get increased; but, the oscillations die after sometime. For the field-field entanglement, it is noticed that the downward fall of the dynamics stops. The atom-field entanglements vanish for this high value of dipole-dipole interaction.
The effects of dipole-dipole interaction for DJCM with SCTS is plotted in Fig. 9. From Fig. 9(a), it can be seen that all the dynamics remain almost unchanged for $g_{d}=0.1$. Only field-field entanglement rises up for large time, making ESDs disappear. For $g_{d}=1.0$, peaks of $C(t)$ come down remarkably; but, the length of ESDs remains same. The field-field entanglement increases more and all the ESDs are removed from the dynamics. The amplitude of $N(t)$ for atom A-field a comes down; however, $N(t)$ for the atom A-field b does not change considerably. For $g_{d}=5$, it is observed that new peaks begin to appear in the dynamics of $C(t)$ and all the ESDs get removed in the dynamics for $g_{d}=10$. The field-field entanglement remains high as before. However, the atom-field entanglements almost disappear for $g_{d}=10$.
6 Effects of spin-spin Ising interaction between the two atoms on entanglement
and
where $J_{z}$ is the coupling strength between the two atoms. $J_{z}$ has the unit of energy. Recently,there are studies on the Ising-type interaction for different configurations. Ghoshal et. al., have investigated the entanglement dynamics of the quenched disordered double Jayens-Cummings model in the presence of Ising-type interaction in the Hamiltonian[31]. In another work, Pandit et. al., have also studied the effects of Ising-type interaction on the entanglement dynamics in the DJCM [28]. In Ref.[33], Laha has analyzed the effects of Ising-type interaction on entanglement in DJCM for vacuum, coherent and thermal states of radiation fields. Sadiek et. al.,[81] have investigated the time evolution and asymptotic behaviour of entanglement with including the dipole-dipole and Ising-type interactions in the Hamiltonian.
The effects of Ising interaction on entanglement for IDDJCM are shown in Fig. 10. The plots show very different effects as are observed in DJCM for SCS and G-L states. From Fig. 10(a) and 10(b), it can be seen that for $J_{z}=0.1$ and $0.3$, we cannot distinguish the plots from the plots with $J_{z}=0$ for $C(t)$ and $N(t)$ (see Figs. 1 and 10). However, for atom A-field b entanglement, the amplitude of $N(t)$ comes down and ESDs are also present in the dynamics. In the case of field a-field b entanglement, for the above mentioned values of $J_{z}$, we observe that the ESDs are removed from the dynamics and amplitude increases. For $J_{z}=0.7$, small peaks appear along with the large peaks in the dynamics of $C(t)$ (red curve) and also the amplitude decreases. For atom A-field a subsystem, the amplitude of $N(t)$ comes down a little and ESDs are also removed in due course of time. In the case of atom A-field b, we do not see any noticeable change in the dynamics. However, for the field a-field b subsystem, the amplitude increases further, and no ESD is observed. For $J_{z}=1.0$, small peaks in the dynamics of $C(t)$ appear earlier in the dynamics and its height increases with time, while the initial peak decreases. As a result, the duration of ESDs gets smaller. For other subsystems, $N(t)$ follows the dynamics similar to that corresponding to $J_{z}=0.7$, viz., with decreased peak height for atom A-field a entanglement, while increased peak height for atom A-field b and field a-field b subsystems.
7 Effects of Kerr-nonlinearity on the entanglement dynamics
@@
The effects of $\chi$ on entanglement dynamics of different subsystems for SCTS in DJCM are presented in Fig. 13. For $\chi=0.1$, ESD in $C(t)$ increases considerably since the smaller peaks disappear from the dynamics (Fig. 13(a), blue curve). For the other subsystems, the dynamics of $N(t)$ does not change considerably for small values of $\chi$. As $\chi$ is increased to $0.3$, the amplitudes of $C(t)$ and $N(t)$ increase; however, ESDs are not removed from the dynamics. For other entanglements, $N(t)$ for atom A-field a and field a-field b decrease prominently with the increasing nonlinearity in the system. However, $N(t)$ for the atom A-field b subsystem increases and ESDs are removed from the dynamics. On further increasing the value of $\chi$ to $0.7$ and $1.0$, $C(t)$ increases very significantly with greater amplitude and ESDs are removed from the dynamics. Similar behaviour is observed for atom A-field b entanglement dynamics. On the other hand, for atom A-field a and field-field entanglement, $N(t)$ decreases prominently with these increased values of $\chi$. This shows that with the addition of $\chi$ to the system, entanglement is transferred in the atom-atom and atom A-field b subsystems from the atom A-field b and field a-field b subsystems.
8 Effects of detuning on the entanglement
@@
To study the effects of detuning $\bar{n}_{c}=2,\bar{n}_{s}=1.0$ and $\bar{n}_{th}=0.1$ are chosen while the values of $\Delta$ are taken to be $2,5,10$. The effects of detuning on entanglements for SCTS in IDDJCM is plotted in Fig. 14. From the plots it is very much noticeable that $\Delta$ plays a significant role in removing ESDs from the system. For $\Delta=2$, length of ESDs get decreased in the dynamics of $C(t)$ (blue curve in Fig. 14(a)), by introducing new additional smaller peaks in the dynamics. Another important observation is that detuning decreases the peak heights of the entanglement pulses significantly with increasing time; however, the periodic nature of $C(t)$ is present to some extent. For atom A-field a entanglement, we see that the addition of $\Delta$ in the system totally destroys the periodic nature of the dynamics of $N(t)$. The amplitude of $N(t)$ is increased and all the ESDs are removed from the dynamics. The periodic-pulse like nature of $N(t)$ for atom A-field b is also gets destroyed by the addition of $\Delta$ to the system. For this smaller value of $\Delta$, the length of ESDs get smaller; however, they are not removed from the dynamics. In the case of field a-field b subsystem, it is observed that the addition of $\Delta=2$ removes all the ESDs from the dynamics of $N(t)$. Although, the periodic nature of $N(t)$ is present in the dynamics, the form of the periodic nature is distorted significantly.
@@
For $\Delta=5$, the peak-heights of $C(t)$ come down further with smaller peaks appearing in some places of the dynamics. This makes the length of ESDs longer at some places and smaller at other places. Also, the entanglement pulses are destroyed with increasing $\Delta$. $N(t)$ for atom A-field a does not change considerably, while for atom A-field b, the amplitude of $N(t)$ increases and some of the ESDs get removed from the dynamics. For field a-field b subsystem, it is observed that the amplitude of $N(t)$ becomes almost half of the initial peak and the periodicity of the dynamics is totally lost.
@@
As $\Delta$ is increased to $\Delta=10$, the length of initial ESD increases as a result of disappearance of other peaks from the dynamics. The dynamics becomes oscillatory for a short while, and again ESDs appear. So, we can see that detuning is not much effective to remove the ESDs from the dynamics of $C(t)$ and increases the value of atom-atom entanglement which is observed for SCS and G-L states.
@@
The effects of $\Delta$ for SCTS in DJCM is depicted in Fig. 15. For $\Delta=2$, initial ESD is removed from the dynamics of $C(t)$; however, the peak following the second long ESD is lost (see black curve in Fig. 3(a) and blue curve in Fig. 15(a)), which makes the second ESD even longer. For all the other subsystems, the addition of $\Delta$ decreases the amplitude of $N(t)$. For the higher values of $\Delta=5,10$, though the amplitude of $C(t)$ increases more initially, the length of subsequent ESDs increases significantly. In the case of other subsystems, $N(t)$ is decreased further which makes ESD appear initially in field-field entanglement. From these plots it is evident that presence of thermal photons forbids entanglement transfer to atom-atom subsystem from other subsystems.
@@ 9 Conclusion
In this work, we have investigated the entanglement dynamics of the atom-atom, atom-field and field-field subsystems with squeezed coherent thermal states in the intensity-dependent double Jaynes-Cummings model (IDDJCM) and the standard double Jaynes-Cummings model (DJCM). It is noticed that the atom-atom and field-field entanglements remain almost unchanged with increasing number of squeezed photons while keeping coherent and thermal photons constant for IDDJCM. An array of entanglement pulses is formed with almost same ESDs and same peak amplitude. However, the atom-field entanglements are affected noticeably. For DJCM, all the dynamics change with increasing squeezed photons and length of ESDs also increases for atom-atom and atom A-field b subsystems.
The addition of photon exchange interaction between the cavities removes ESDs from field-field entanglement dynamics and from the atom-atom for large values of $\kappa$, for both IDDJCM and DJCM. The atom-field entanglements vanish for large values of $\kappa$. In the case of DJCM, formation of wave packets is observed for field-field subsystem.
Like photon exchange interaction, dipole-dipole interaction removes ESDs from the dynamics of $C(t)$ and $N(t)$ of field-field subsystem for high values of $g_{d}$ for both IDDJCM and DJCM. However, in this case, formation of wave packets is not observed. Ising interaction is effective to remove ESDs from field-field entanglement for IDDJCM; however, it is unable to remove ESDs from atom-atom and atom-field entanglements. For DJCM, the dynamics of $C(t)$ and $N(t)$ are not affected considerably by the addition of Ising interaction in the system.
Kerr-nonlinearity effectively removes ESDs from the all the entanglement dynamics. For IDDJCM, $C(t)$ and $N(t)$ for all the subsystems increase for higher values of $\chi$. In the case of DJCM also, $C(t)$ and $N(t)$ for atom-field subsystems increase for higher values of $\chi$; however, field-field entanglement decreases after a critical value of $\chi$. Detuning tries to remove ESDs from all the subsystems. ESDs vanish from field-field entanglement; however its peak-amplitude comes down with increasing value of $\Delta$. Also, for higher values of detuning, the $C(t)$ increases more and other entanglements decrease, which is an implications of entanglement transfer from other subsystems to atom-atom subsystem.